Put the equation of a line into slope-intercept form. 3x + 2y = 10
Answer by nerdybill(7384)

Put the equation of a line into slope-intercept form. 3x + 2y = 10
.
Starting with:
3x + 2y = 10
Subtract 3x from both sides:
2y = -3x + 10
Divide both sides by 2:
y = (-3/2)x + 5 (this is what they're looking for)
.
slope-intercept form:
y = mx + b
where
m is the slope
b is the y-intercept